OpenStudy (anonymous):

The height in inches (h) of a tomato plant growing in a garden is given by the equation h=2t+3, where t represents time in days. When will the tomato plant be between 13 and 27 inches?

OpenStudy (anonymous):

When will the tomato plant be between 13 and 27 inches, means you need to substitute h=13 and 27 to fins the two values of t and that value of t will give you the ans since h=2t+3 i.e. at h=13 we find 13= 2t+3 i.e. 13-3= 2t 10=2t i.e. t=10/2 i.e. t=5 days similarly at h=27 we find 27= 2t+3 i.e. 27-3= 2t 24=2t i.e. t=24/2 i.e. t=12 days hence in 5 and 12 days the height of the tomato plant will be between 13 and 27 inches respectively.

OpenStudy (anonymous):

\[13 \le 2t+3 \le 27\] \[13-3\le 2t \le 27-3 \implies 10 \le 2t \le 24 \] \[\implies 5 \le t \le 12\]